Two travelling waves y1(x,t) = A sin⁡[k(x-vt)] and y2(x,t) = A sin[k(x+vt)] are superimposed on the same string. The distance between two adjacent nodes is:

1
Expert's answer
2021-12-20T10:29:34-0500

"y=y_1+y_2=2A\cos(vt)\sin(kx),"

"d=\frac{\pi}k."